# Simultaneous Multi-Robot Motion Planning with Projected Diffusion Models

### v1

@@ 4.2 Augmented Lagrangian for Efficient Projections
The Lagrangian dual function of the MRMP problem is thus defined as:
@@
where ${\cal L}(\bm{\Pi})$ represents the original MRMP objective function, $\boldsymbol{\nu}_{a}$ and $\boldsymbol{\nu}_{o}$ are Lagrangian multipliers, and $\mathcal{H}_{a}$ and $\mathcal{H}_{o}$ represent the equality constraints enforcing inter-robot collision avoidance and obstacle avoidance, respectively. While the standard Lagrangian formulation provides a means of incorporating constraints into the optimization process, it suffers from slow convergence due to the instability of dual variable updates, particularly in nonconvex settings. To mitigate these issues, we adopt the augmented Lagrangian method (Boyd et al., 2011; Kotary et al., 2022), which introduces additional penalty terms on the constraint residuals:
@@
At each iteration, the primal variable $\bm{\Pi}^{k}$ is updated by minimizing the augmented Lagrangian, and the dual variables $\bm{\nu}_{a},\bm{\nu}_{o}$ are updated via a gradient ascent step on the constraint residuals. As the paper will show next, this approach is key to enabling the generation of collision-free and kinematic viable trajectories for multiple robots, especially in complex scenarios. The augmented projection process is described in Algorithm 2.
@@
1:  Input: Tolerance $\delta$, Weight $\rho$, Initial Trajectory $\hat{\boldsymbol{\Pi}}$
2:  while $\nabla_{\boldsymbol{\nu}_{a}}<\delta\ \textbf{and}\ \nabla_{\boldsymbol{\nu}_{o}}<\delta$ do
@@
6:     $\rho\leftarrow\text{Update}(\rho)$
@@ 5 Experimental Settings
Maps. We evaluate MRMP algorithms on both randomly generated and structured, real-world-inspired maps.
$\bullet$ Random maps include environments with increasing levels of complexity: empty maps test inter-robot collision avoidance in open spaces; basic maps introduce 10 obstacles, requiring navigation of narrow corridors; dense maps contain 20 obstacles, significantly restricting movement and increasing planning difficulty. The progression from empty to dense maps allows us to systematically assess the algorithms on spatial constraints and coordination.
$\bullet$ Practical maps simulate real-world environments such as warehouses and buildings, where constrained pathways and structured layouts impose additional planning challenges. Corridor maps replicate narrow passages where robots must coordinate movements to pass each other without deadlocks. Shelf maps mirror warehouse storage layouts with tight aisles, requiring precise navigation. Room maps introduce multiple rooms connected by doors, estricting the number of robots that can enter at the same time and requiring careful scheduling to prevent congestion. Compared to random maps, practical maps emphasize not only collision avoidance but also global coordination, as robots must find feasible routes through constrained spaces.
Each scenario contains 25 distinct maps.
@@
Task assignment. Start and goal states are assigned differently for random and practical maps. In random maps, they are placed at feasible positions without collisions (see Figure 2), while in practical maps, they are constrained to predefined zones that reflect real-world constraints, such as pickup and drop-off locations in warehouses (see Figure 3). To ensure that each instance has feasible solutions, we discretize the environment and use multi-agent path finding algorithms to verify whether a collision-free solution exists. If no feasible assignment is found, the configuration is regenerated. We conduct experiments with 3, 6, and 9 robots, generating 10 test cases for each configuration, except for corridor maps, where we use 2 robots.
@@
    Diffusion Models (DM): We adopt standard diffusion models fine-tuned over the benchmark trajectories to address MRMP directly (Nichol & Dhariwal, 2021).
@@
    Multi-robot Motion Planning Diffusion (MMD): A recently introduced solution that combines diffusion models with classical search-based techniques—generating MRMP solutions under collision constraints (Shaoul et al., 2024).
We also compare our proposed approaches with classical algorithms in Appendix B.
@@ 6 Experiments
In this section, we evaluate the performance of various MRMP algorithms using our proposed benchmark. In most cases, regular projection methods cannot be applied to obtain solutions in MRMP due to the increased number of robots and obstacles (hence, our omission of Christopher et al. (2024) as a baseline). Thus, for our implementation of SMD, we use the projection with the augmented Lagrangian method for our experiments, except for the narrow-corridor map. The implementation details can be found in Appendix A. Additionally, we introduce supplementary details in the appendix B, including the results of classical algorithms and additional metrics.
@@
6.1 Performance on Random Maps
First, notice that the Standard Diffusion Model is unviable for MRMP tasks. Despite reporting a reasonable success rate (78%) for 3 robots in empty maps, when simple obstacles are introduced, the success rate falls dramatically (5%). Additionally, DM fails to produce any feasible trajectory when scaling to six or nine robots, even in empty maps.
Next, we focus on Motion Planning Diffusion. While it reports near-perfect success rates for the 3-robots setting (the smallest in our benchmark), this approach is highly ineffective when scaled to additional robots. In empty environments, MPD reports 35% success rates and it fails to synthesize any feasible trajectory when obstacles are introduced to the map or when scaling to nine robots.
Multi-robot Motion Planning Diffusion represents the current state-of-the-art method for MRMP. This method outperforms other baselines in its ability to handle an increased number of robots. For instance, it reports non-zero success rates for up to nine robots. However, there remains a steep performance drop with the increasing number of robots. For instance, it reports a success rate of 27% in dense maps for 9 agents. A similar decrease in constraint satisfaction occurs for in other scenarios besides the empty maps.
In contrast, Simultaneous MRMP Diffusion provides new state-of-the-art results for both constraint satisfaction and path length. SMD reports perfect success rates and collision ratios for all tasks except the most complex maps (dense environments) and 9 robots, and, even in this case, it provides near-perfect results (96% successful). This is a 3.6x improvement over the previous state-of-the-art. Sample trajectories generated by SMD are visualized in Figure 5.
@@ 6.2 Performance on Practical Maps
Figure 6: Results for each method on corridor maps with three different numbers of robots. Gray bars represents the failure rate, and values on top of the bars indicate average path length per robot.
@@ 7 Discussion and Limitations
While SMD provides state-of-the-art results for MRMP tasks, forethought should be given as to the appropriate applications of this approach. First, generative models incur inference-time costs that should be considered when applying these models to motion planning. Diffusion models are most advantageous in scenarios where the complexity of the environment and the need for multimodal trajectory generation outweigh the benefits of faster but more rigid methods. For instance, such architectures are ideal for handling unstructured environments with dense obstacles, dynamic changes, and uncertain sensor inputs, where traditional methods struggle to adapt but may be unnecessary in structured or open spaces.
@@ 8 Conclusion
This work introduced Simultaneous MRMP Diffusion (SMD), a method that integrates constrained optimization with generative diffusion models to generate collision-free, kinematically feasible trajectories for multi-robot systems. By alternating diffusion sampling with projections onto the feasibility set, SMD eliminates the need for rejection sampling or post-processing. To handle complex nonconvex constraints, the paper incorporates an Augmented Lagrangian Method enabling scalability to challenging multi-robot scenarios where other methods fail. To support rigorous evaluation, a benchmark suite was also introduced, covering environments that reflect real-world motion planning challenges. Extensive experiments across varying obstacle densities and increasing robot counts demonstrated that SMD achieves significantly higher success rates and better objective values than competing methods. The results highlight its robustness in handling both static obstacles and dynamic interactions, making it a promising approach for real-world MRMP applications.
Future work will focus on improving scalability through decomposition techniques and constraint relaxations.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.3 Evaluation Details
Appendix B Additional Experimental Results
@@
As shown in Table 3 and Table 4, our proposed SMD consistently achieves an impressive success rate across all tested scenarios. In contrast, traditional approaches such as EECBS also maintain a high success rate but struggle with path length, which is approximately 20% longer than that of our method. This is caused by they operate within a predefined grid structure. Notably, our SMD achieves near-zero collision rates, outperforming other learning-based approaches, which frequently fail to avoid collisions in complex scenarios.
@@
Map Robots Metric EECBS DM MPD MMD Ours Corridor Maps 2 S $\uparrow$ 0.7120 0 0 0 1 L $\downarrow$ 1.4208 1.6619 A $\downarrow$ 0.0012 C $\downarrow$ 1 1 1 0 Shelf Maps 3 S $\uparrow$ 1 0.1200 0.5960 0.5960 1 L $\downarrow$ 1.1697 0.7884 0.8825 0.9371 0.9017 A $\downarrow$ 0.0030 0.0081 0.0030 0.0060 C $\downarrow$ 0.5027 0.2493 0.3133 0 6 S $\uparrow$ 1 0 0 0.5000 0.9560 L $\downarrow$ 1.1493 0.9317 1.0632 A $\downarrow$ 0.0030 0.0018 C $\downarrow$ 1 0.9960 0.3740 0.01 9 S $\uparrow$ 1 0 0 0.4480 0.9320 L $\downarrow$ 1.1427 0.9331 1.0838 C $\downarrow$ 0.0032 0.0068 A $\downarrow$ 1 1 0.4218 0.0111 Room Maps 3 S $\uparrow$ 1 0 0.0840 0.1360 1 L $\downarrow$ 1.6956 1.1771 1.2697 1.1917 A $\downarrow$ 0 0.0066 0.0043 0.0030 C $\downarrow$ 1 0.6520 0.6867 0 6 S $\uparrow$ 0.9960 0 0 0.024 1 L $\downarrow$ 1.6778 1.2660 1.3019 A $\downarrow$ 0.0032 0.0014 C $\downarrow$ 1 1 0.8033 0 9 S $\uparrow$ 0.9920 0 0 0 0.9600 L $\downarrow$ 1.6620 1.3374 A $\downarrow$ 0.0033 C $\downarrow$ 1 1 0.8849 0.0044

### v2

@@ 4.3 Lagrangian Relaxation for Efficient Projections
The Lagrangian function of the MRMP problem is thus defined as:
@@
where ${\cal J}(\bm{\Pi})$ represents the objective function, which typically finds the nearest feasible point to the input in the projection, as shown in Eq. 4. $\boldsymbol{\nu}_{a}$ and $\boldsymbol{\nu}_{o}$ are Lagrangian multipliers, and $\mathcal{H}_{a}$ and $\mathcal{H}_{o}$ represent the equality constraints enforcing inter-robot collision avoidance and obstacle avoidance, respectively. While the standard Lagrangian formulation provides a means of incorporating constraints into the optimization process, it suffers from slow convergence due to the instability of dual variable updates, particularly in nonconvex settings. To mitigate these issues, we adopt the augmented Lagrangian method (Boyd et al., 2011; Kotary et al., 2022) to improve the convergence performance, which introduces additional penalty terms on the constraint residuals:
@@
At each iteration, the primal variable $\bm{\Pi}^{k}$ is updated by minimizing the augmented Lagrangian, and the dual variables $\bm{\nu}_{a},\bm{\nu}_{o}$ are updated via a gradient ascent step on the constraint residuals. This approach enables the generation of collision-free and kinematic viable trajectories for multiple robots, especially in complex scenarios. The efficient projection process is described in Algorithm 2.
@@
1:  Input: Tolerance $\delta_{a}$ and $\delta_{o}$, Weight $\rho$, Initial Trajectory $\hat{\boldsymbol{\Pi}}$, scaling factor $\zeta$
2:  while $\nabla_{\boldsymbol{\nu}_{a}}<\delta_{a}\ \textbf{and}\ \nabla_{\boldsymbol{\nu}_{o}}<\delta_{o}$ do
@@
6:     $\rho\leftarrow\zeta\times\rho$
@@
4.4 Theoretical Analysis for SMD
Building on our earlier discussion of how the repeated projection mechanism ensures convex constraint satisfaction, we now present a more detailed theoretical justification of our proposed SMD for MRMP. The key insight is SMD can provide feasible trajectories for MRMP by introducing the Lagrangian relaxation method. Although the full MRMP problem is nonconvex, we leverage the constraint structure to control violations during each projection step.
For simplicity, we conduct our analysis on the standard Lagrangian formulation in Eq. (9), as the augmented Lagrangian method mainly enhances the convergence of the original Lagrangian relaxation. Before proceeding, we make the following assumptions commonly satisfied in MRMP formulations.
Assumption 4.1.
The cost function of relaxed MRMP $\mathcal{L}(\boldsymbol{\Pi},\boldsymbol{\nu}_{a},\boldsymbol{\nu}_{o})$ is continuously differentiable and convex over the convex set $\Omega_{c}$.
Proposition 4.2 (Convex Feasibility Guarantee).
Let ${\bm{\Pi}}$ be the trajectory output by the projection operator $\mathcal{P}_{\Omega}(\bm{\Pi})$ within the diffusion step of SMD. Suppose that Assumption 4.1 holds. Then, the generated trajectory ${\bm{\Pi}}$ satisfies
$\text{dist}(\boldsymbol{\Pi},\Omega_{c})\leq\xi$
where $\text{dist}(\cdot)$ denotes the distance between the trajectory $\boldsymbol{\Pi}$ and the feasible region $\Omega_{c}$, and $\xi\geq 0$ can be made arbitrarily small.
Remark 4.3.
This proposition provides a theoretical guarantee that our SMD can ensure the feasibility for convex feasible set $\Omega_{c}$ by introducing Lagrangian relaxation methods. In addition to the convex constraints, SMD handles nonconvex constraints defined by functions $\mathcal{H}_{a}(\bm{\Pi})$ and $\mathcal{H}_{o}(\bm{\Pi})$ via a dual ascent method with the user-defined stopping criterion $\delta_{a}$ and $\delta_{o}$. Specifically, SMD ensures the output of the projection operator ${\bm{\Pi}}$:
$||\mathcal{H}_{a}({\bm{\Pi}})||\leq\delta_{a},\quad||\mathcal{H}_{o}({\bm{\Pi}})||\leq\delta_{o},$
for some tolerances $\delta_{a},\delta_{o}>0$.
@@ 5 Experimental Settings
Maps. We evaluate MRMP algorithms on both randomly generated and real-world-inspired maps.
$\bullet$ Random maps include environments with increasing levels of complexity: empty maps test inter-robot collision avoidance in open spaces; basic maps introduce 10 obstacles, requiring navigation within the feasible region that excludes obstacle-occupied areas ; dense maps contain 20 obstacles, significantly restricting movement and increasing planning difficulty. The progression from empty to dense maps allows us to systematically assess the algorithms on spatial constraints and coordination.
$\bullet$ Practical maps simulate real-world environments such as warehouses and buildings, where constrained pathways and structured layouts impose additional planning challenges. Corridor maps replicate narrow passages where robots must coordinate movements to pass each other without deadlocks. Shelf maps mirror warehouse storage layouts with tight aisles, requiring precise navigation. Room maps introduce multiple rooms connected by doors, restricting the number of robots that can enter at the same time and requiring careful scheduling to prevent congestion. Compared to random maps, practical maps emphasize not only collision avoidance but also global coordination, as robots must find feasible routes through constrained spaces.
Each scenario includes 25 maps with different obstacle configurations.
@@
Task assignment. Start and goal states are assigned differently for random and practical maps. In random maps, they are placed at feasible positions without collisions (see Figure 2), while in practical maps, they are constrained to predefined zones that reflect real-world constraints, such as pickup and drop-off locations in warehouses (see Figure 3). We conduct experiments with 3, 6, and 9 robots, generating 10 test cases for each configuration, except for corridor maps, where we use 2 robots.
@@
    Diffusion Models (DM): We adopt standard diffusion models trained over the feasible trajectories to address MRMP directly (Nichol & Dhariwal, 2021).
@@
    Multi-robot Motion Planning Diffusion (MMD): A recently introduced solution that combines diffusion models with classical search-based techniques—generating MRMP solutions under collision constraints (Shaoul et al., 2025).
We also compare our proposed approaches with classical algorithms in Appendix C.
@@ 6 Experiments
In this section, we evaluate the performance of various MRMP algorithms using our proposed benchmark. In most cases, regular projection methods cannot be applied to obtain solutions in MRMP due to the huge computational burden (hence, our omission of Christopher et al. (2024) as a baseline). Thus, for our implementation of SMD, we use the projection with the Lagrangian relaxation method for our experiments, except for the narrow-corridor map. The implementation details can be found in Appendix A. Additional experimental results are presented in Appendix C, including the performance of classical algorithms, additional evaluation metrics, runtime comparisons, sensitivity analysis of projection parameters, and results on other map types.
@@
6.1 Performance on Random Maps
First, notice that the Standard Diffusion Model (DM) is unviable for MRMP tasks. Despite reporting a reasonable success rate (78%) for 3 robots in empty maps, when simple obstacles are introduced, the success rate falls dramatically (5%). Additionally, DM fails to produce any feasible trajectory when scaling to six or nine robots, even in empty maps.
Next, we focus on Motion Planning Diffusion (MPD). While it reports near-perfect success rates for the 3-robots setting (the smallest in our benchmark), this approach is highly ineffective when scaled to additional robots. In empty environments, MPD reports 35% success rates with six robots and it fails to synthesize any feasible trajectory when obstacles are introduced to the map or when scaling to nine robots.
Multi-robot Motion Planning Diffusion (MMD) represents the current state-of-the-art method for MRMP. This method outperforms other baselines in its ability to handle an increased number of robots. For instance, it reports non-zero success rates for up to nine robots. However, there remains a steep performance drop with the increasing number of robots. For instance, it reports a success rate of 27% in dense maps for 9 agents. A similar decrease in constraint satisfaction occurs for in other scenarios besides the empty maps.
In contrast, Simultaneous MRMP Diffusion (SMD) provides new state-of-the-art results for both constraint satisfaction and path length. Unlike other methods whose success rates rapidly decline as robot and obstacle numbers grow, SMD maintains feasibility in scenarios with increased robot and obstacle counts (e.g., dense maps with 9 robots and 20 obstacles). Specifically, SMD is the only known method that provides feasible solutions for the largest number of robots in complex environments. SMD reports perfect success rates and collision ratios for all tasks except the most complex maps (dense environments) and 9 robots. Even in the most challenging dense maps, where it still provides near-perfect results (96% successful). This is a 3.6x improvement over the previous state-of-the-art. Sample trajectories generated by SMD are visualized in Figure 5.
@@ 6.2 Performance on Practical Maps
Figure 6: Results for each method on corridor maps. Gray bars represents the failure rate, and values on top of the bars indicate average path length per robot.
@@ 7 Discussion and Limitations
While SMD provides state-of-the-art results for MRMP tasks, consideration should be given to the appropriate applications of this approach. First, generative models incur inference-time costs that should be considered when applying these models to motion planning. Diffusion models are most advantageous in scenarios where the complexity of the environment and the need for multimodal trajectory generation outweigh the benefits of faster but more rigid methods. For instance, such architectures are ideal for handling unstructured environments with dense obstacles, dynamic changes, and uncertain sensor inputs, where traditional methods struggle to adapt but may be unnecessary in structured or open spaces.
@@ 8 Conclusion
This work introduced Simultaneous MRMP Diffusion (SMD), a method that integrates constrained optimization with generative diffusion models to generate collision-free, kinematically feasible trajectories for multi-robot systems. By alternating diffusion sampling with projections onto the feasibility set, SMD eliminates the need for rejection sampling or post-processing. To handle complex nonconvex constraints, the paper incorporates an Augmented Lagrangian Method enabling scalability to challenging multi-robot scenarios where other methods fail. To support rigorous evaluation, a benchmark was also introduced, covering environments that reflect real-world motion planning challenges. Extensive experiments across varying obstacle densities and increasing robot counts demonstrated that SMD achieves significantly higher success rates and better objective values than competing methods. The results highlight its robustness in handling both static obstacles and dynamic interactions, making it a promising approach for real-world applications.
There are several promising directions for future work. Extending SMD to more complex scenarios, such as 3D environments and humanoid locomotion is very promising. Furthermore, the extension of SMD to decentralized settings still remain to be explored.
Acknowledgments
This research is partially supported by NSF grants 2334936, 2334448, and NSF CAREER Award 2401285. The research at the University of California, Irvine was supported by the NSF grants 2434916, 2346058, 2321786, 2121028, and 1935712, as well as gifts from Amazon Robotics. The authors acknowledge Research Computing at the University of Virginia for providing computational resources that have contributed to the results reported within this paper. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government.
@@ Acknowledgments
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.3 Evaluation Details
Appendix B Missing Proofs
B.1 Proof of Proposition 4.2
Proof.
By Assumption 4.1, the projection operator $\mathcal{P}_{\Omega}(\bm{x})$ performs the following optimization problem at each iteration:
$\boldsymbol{\Pi}=$ $\arg\min_{\boldsymbol{\Pi}\in\Omega_{c}}\ \mathcal{L}(\boldsymbol{\Pi},\boldsymbol{\nu}_{a},\boldsymbol{\nu}_{o}).$ (13)
Since $\Omega_{c}$ is convex and $\mathcal{L}(\boldsymbol{\Pi},\boldsymbol{\nu}_{a}^{k},\boldsymbol{\nu}_{o}^{k})$ is convex and continuously differentiable over $\Omega_{c}$, the optimization problem has a unique global minimum.
We define a single update step for the diffusion sampling process as:
$\mathcal{U}(\boldsymbol{\Pi}_{t}^{i})=\boldsymbol{\Pi}_{t}^{i}+\gamma_{t}\mathbf{s}_{\theta}(\boldsymbol{\Pi}_{t}^{i},t)+\sqrt{2\gamma_{t}}\mathbf{z}.$ (14)
For each time step $t$, there exists a minimum iteration index $i=\bar{I}$ such that:
$\exists{\bar{I}}\;\texttt{s.t.}\;\left\|(\boldsymbol{\Pi}_{t}^{\bar{I}}+\gamma_{t}\nabla_{\boldsymbol{\Pi}_{t}^{i}}\log q(\boldsymbol{\Pi}_{t}|\boldsymbol{\Pi}_{0}))\right\|_{2}\leq\left\|F_{t}\right\|_{2}$ (15)
where $F_{t}$ is the closest point to the global optimum that can be reached via a single gradient step from any point in $\Omega_{c}$.
Let $Error$ be the distance between $\boldsymbol{\Pi}_{t}^{i}$ and its nearest feasible point. Using Theorem 5.2 in (Christopher et al., 2024), for any $i\geq I$, we have:
$\mathbb{E}\left[\textit{Error}(\mathcal{U}(\mathcal{P}_{\Omega_{c}}(\boldsymbol{\Pi}_{t}^{i})),\Omega_{c})\right]\leq\xi\leq\mathbb{E}\left[\textit{Error}(\mathcal{U}(\boldsymbol{\Pi}_{t}^{i}),\Omega_{c})\right].$ (16)
where $\xi\geq 0$ is arbitrarily small.
Therefore, the expected distance between the final generated trajectory and the feasible set $\Omega_{c}$ is bounded above by $\xi$, which means our SMD method yields a strictly smaller expected constraint violation and provides a feasibility guarantee for the convex constraint set $\Omega_{c}$.
Appendix C Additional Experimental Results
@@
As shown in Table 3 and Table 4, our proposed SMD consistently achieves an impressive success rate across all tested scenarios. In contrast, traditional approaches such as EECBS also maintain a high success rate but struggle with path length, which is approximately 20% longer than that of our method. This is caused by they operate within a predefined grid structure. Notably, our SMD achieves near-zero collision rates, outperforming other learning-based approaches, which frequently fail to avoid collisions in complex scenarios. The computational costs are reported in Table 5. Notably, more challenging tasks typically take longer to solve, thus increasing the average running time.
In addition, Table 6 shows evaluation results on practical maps introduced by (Shaoul et al., 2025). SMD still maintains a zero collision rate. These results further confirm the practical applicability and reliability of our approach.
In Figures 9 and 10, we present a sensitivity analysis of the scaling factor $\alpha$ used in our ALM-based projection. The results suggest that careful tuning of $\zeta$ is not necessary; using a moderate value such as 1.05 already achieves good convergence performance.
@@
Map Robots Metric EECBS DM MPD MMD Ours Corridor Maps 2 S $\uparrow$ 1 0 0 0 1 L $\downarrow$ 1.4208 1.6619 A $\downarrow$ 0.0012 C $\downarrow$ 1 1 1 0 Shelf Maps 3 S $\uparrow$ 1 0.1200 0.5960 0.5960 1 L $\downarrow$ 1.1697 0.7884 0.8825 0.9371 0.9017 A $\downarrow$ 0.0030 0.0081 0.0030 0.0060 C $\downarrow$ 0.5027 0.2493 0.3133 0 6 S $\uparrow$ 1 0 0 0.5000 0.9560 L $\downarrow$ 1.1493 0.9317 1.0632 A $\downarrow$ 0.0030 0.0018 C $\downarrow$ 1 0.9960 0.3740 0.01 9 S $\uparrow$ 1 0 0 0.4480 0.9320 L $\downarrow$ 1.1427 0.9331 1.0838 C $\downarrow$ 0.0032 0.0068 A $\downarrow$ 1 1 0.4218 0.0111 Room Maps 3 S $\uparrow$ 1 0 0.0840 0.1360 1 L $\downarrow$ 1.6956 1.1771 1.2697 1.1917 A $\downarrow$ 0 0.0066 0.0043 0.0030 C $\downarrow$ 1 0.6520 0.6867 0 6 S $\uparrow$ 0.9960 0 0 0.0240 1 L $\downarrow$ 1.6778 1.2660 1.3019 A $\downarrow$ 0.0032 0.0014 C $\downarrow$ 1 1 0.8033 0 9 S $\uparrow$ 0.9920 0 0 0 0.9600 L $\downarrow$ 1.6620 1.3374 A $\downarrow$ 0.0033 C $\downarrow$ 1 1 0.8849 0.0044
Table 5: Running time in seconds with success rates (shown in parentheses and expressed as percentages) for all learning-based methods across all maps. We consider only the time consumed to generate feasible trajectories for MMD, as the running time for failed trajectories in MMD is determined by a predefined time limit, which could be meaningless if we calculate the running time using time limit (e.g., 3000s).
Map Robots DM MPD MMD Ours Empty Maps 3 4.2(78.4) 14.1(100) 37.6(100) 29.5(100) 6 3.7(0) 11.7(35.2) 72.1(100) 103.8(100) 9 3.9(0) 9.9(0) 108.5(100) 204.7(100) Basic Maps 3 4.1(5.2) 14.3(99.6) 39.6(92.8) 75.1(100) 6 4.1(0) 9.9(0) 81.9(83.2) 235.2(100) 9 4.1(0) 10.2(0) 123.1(72.8) 481.6(100) Dense Maps 3 3.8(0) 14.5(99.2) 46.3(66.0) 73.7(100) 6 4.3(0) 10.1(0) 105.3(43.6) 255.1(100) 9 4.1(0) 10.3(0) 192.4(27.2) 551.1(96.0) Shelf Maps 3 4.1(12.0) 9.3(59.6) 59.2(59.6) 91.3(100) 6 4.1(0) 7.6(0) 128.4(50.0) 283.2(95.6) 9 4.1(0) 8.2(0) 200.2(44.8) 585.9(93.2) Room Maps 3 4.1(0) 7.9(8.4) 20.5(13.6) 74.6(100) 6 4.1(0) 8.8(0) 41.2(2.4) 158.7(100) 9 4.1(0) 8.5(0) N/A(0) 255.7(96.0)
Table 6: Additional results for MMD and SMD in maps used in (Shaoul et al., 2025).
Map Robots Metric MMD Ours Highways Maps 3 S $\uparrow$ 1 1 L $\downarrow$ 1.1638 1.1391 A $\downarrow$ 0.0026 0.0030 C $\downarrow$ 0 0 6 S $\uparrow$ 1 1 L $\downarrow$ 1.1669 1.1775 A $\downarrow$ 0.0028 0.0022 C $\downarrow$ 0 0 9 S $\uparrow$ 1 1 L $\downarrow$ 1.1792 1.1802 A $\downarrow$ 0.0028 0.0031 C $\downarrow$ 0 0 Conveyor Maps 3 S $\uparrow$ 1 1 L $\downarrow$ 1.1529 1.1499 A $\downarrow$ 0.0030 0.0029 C $\downarrow$ 0 0 6 S $\uparrow$ 1 1 L $\downarrow$ 1.1726 1.1806 A $\downarrow$ 0.0027 0.0044 C $\downarrow$ 0 0 9 S $\uparrow$ 1 1 L $\downarrow$ 1.1916 1.2003 A $\downarrow$ 0.0031 0.0067 C $\downarrow$ 0 0 Drop-Region Maps 3 S $\uparrow$ 1 1 L $\downarrow$ 1.1417 1.0337 A $\downarrow$ 0.0024 0.0051 C $\downarrow$ 0 0 6 S $\uparrow$ 1 1 L $\downarrow$ 1.1604 1.1841 A $\downarrow$ 0.0028 0.0039 C $\downarrow$ 0 0 9 S $\uparrow$ 1 1 L $\downarrow$ 1.1719 1.3556 A $\downarrow$ 0.0027 0.0075 C $\downarrow$ 0 0
(a) Empty Maps with 3 Robots.
(b) Empty Maps with 6 Robots.
(c) Empty Maps with 9 Robots.
(d) Basic Maps with 3 Robots.
(e) Basic Maps with 6 Robots.
(f) Basic Maps with 9 Robots.
(g) Dense Maps with 3 Robots.
(h) Dense Maps with 6 Robots.
(i) Dense Maps with 9 Robots.
(j) Shelf Maps with 3 Robots.
(k) Shelf Maps with 6 Robots.
(l) Shelf Maps with 9 Robots.
(m) Room Maps with 3 Robots.
(n) Room Maps with 6 Robots.
(o) Room Maps with 9 Robots.
Figure 9: Sensitivity analysis of the scaling factor $\zeta$ on gradient convergence for inter-agent collision avoidance constraints in ALM-based projection, evaluated for each map with different numbers of robots. The scaling factor $\zeta$ determines how the coefficient of the augmented term is multiplied at each iteration. We test five increasing values of $\zeta$ from 1.01 to 1.09, as well as a constant value of 1.00 as the ablation study.
(a) Basic Maps with 3 Robots.
(b) Basic Maps with 6 Robots.
(c) Basic Maps with 9 Robots.
(d) Dense Maps with 3 Robots.
(e) Dense Maps with 6 Robots.
(f) Dense Maps with 9 Robots.
(g) Shelf Maps with 3 Robots.
(h) Shelf Maps with 6 Robots.
(i) Shelf Maps with 9 Robots.
(j) Room Maps with 3 Robots.
(k) Room Maps with 6 Robots.
(l) Room Maps with 9 Robots.
Figure 10: Sensitivity analysis of the scaling factor on gradient convergence for obstacle collision avoidance constraints in ALM-based projection, which leads to similar conclusions with inter-agent collision avoidance constraints.